Eric has another get-rich-quick idea, but needs funding to support it. He chooses an all-debt funding scenario. He will borrow $4740 from Wendy, who will charge him 8% on the loan. He will also borrow $3150 from Bebe, who will charge him 10% on the loan, and $2110 from Shelly, who will charge him 16% on the loan. What is the weighted average cost of capital for Eric? What is the weighted average cost of capital for Eric?
Answer #1
Total Capital value = Value of Wendy + Value of Bebe + Value of Shelly
=4740+3150+2110
=10000
Weight of Wendy = Value of Wendy/Total Capital Value
= 4740/10000
=0.474
Weight of Bebe = Value of Bebe/Total Capital Value
= 3150/10000
=0.315
Weight of Shelly = Value of Shelly/Total Capital Value
= 2110/10000
=0.211
Cost of Capital = Weight of Wendy*Cost of Wendy+Weight of Bebe*Cost of Bebe+Weight of Shelly*Cost of Shelly
Cost of Capital = 8*0.474+10*0.315+16*0.211
Cost of Capital = 10.318